OpenStudy (anonymous):

find the value of e^-3.4

OpenStudy (anonymous):

\[e{}^{\wedge}-3.4=e{}^{\wedge}(-34/10)=\frac{1}{e^{17/5}}=0.0333733 \]

OpenStudy (anonymous):

I will look at it real quick to make sure I get it!

OpenStudy (anonymous):

Thank you so much!! God bless<><